OpenStudy (anonymous):

You are playing a game in which a single die is rolled. If a 2 or 5 comes up you win $36. Otherwise, you lose $36. What is your expected value for the game?

OpenStudy (anonymous):

\[\text{Expected} \text{ Value}=\left(36*\frac{2}{6}\right)+\left(-36*\frac{4}{6}\right)=- \text{$\$$12} \]